Documenta Mathematica

Band 17, 2012

Markus Reineke Degenerate Cohomological Hall Algebra and Quantized Donaldson-Thomas Invariants for

m-Loop Quivers  1–22

P. Broussous, V. Scherre, and S. Stevens Smooth Representations of GL_{m}(D)

V: Endo-Classes  23–77

Oliver Goertsches The Equivariant Cohomology of Isotropy Actions

on Symmetric Spaces  79–94

Amalendu Krishna

Equivariant Cobordism of Schemes  95–134

S. Baek, E. Neher, K. Zainoulline Basic Polynomial Invariants, Fundamental Representations and the

Chern Class Map  135–150

iv